OpenStudy (anonymous):

Help please. Use the method of substitution to solve the system of liner equations. 2x-y=3 7x-3y=14 The solution is? Type an ordered pair.

OpenStudy (anonymous):

solve the first equation for y,\[y=2x-3\] Now that you know what y is, substitute that into equation 2 and solve for x. Once you have x plug it into equation 1 and obtain y. Do you think you can do this from this explanation?

OpenStudy (anonymous):

I really need help in the steps as I am not good at these equations at all. If you have time to help me. please.

OpenStudy (anonymous):

sure lets start at the beginning. Solve the first equation for one of the variables. I chose to solve for y because it doesn't have a coefficient and it will be less messy.

OpenStudy (anonymous):

add y to both sides of the equation.\[y+2x-y=3+y\]You can see that the y's will cancel and you will be left with \[2x=y+3\]

OpenStudy (anonymous):

next subtract e from both sides \[2x-3=y+3-3; y=2x-3\]

OpenStudy (anonymous):

remember if you do the same thing to each side of the equation essentially nothing changes

OpenStudy (anonymous):

Now that we know y plug it into the 2nd equation.\[7x-3(2x-3)=14\]

OpenStudy (anonymous):

(still listening and taking notes as you go along) :)

OpenStudy (anonymous):

I prefer to write everything using addition and negative numbers. It helps me to not make mistakes in sign changes. \[7x+(-3)(2x+(-3))\]distribute the -3. \[7x+[(-3\times2x)+(-3\times-3)]=14\]

OpenStudy (anonymous):

Now we should have \[7x+(-6x)+9=14\]We can simplify further to \[x+9=14\]

OpenStudy (anonymous):

what do you think we do next to solve for x??

OpenStudy (anonymous):

add 14 to both sides?

OpenStudy (anonymous):

Keep in mind that our goal is to get x by itself on one side of the equation.

OpenStudy (anonymous):

try again...

OpenStudy (anonymous):

9 to both sides?

OpenStudy (anonymous):

add or subtract??

OpenStudy (anonymous):

subtract

OpenStudy (anonymous):

Correcttt!!!!!!!

OpenStudy (anonymous):

so we should arrive at \[x=5\]

OpenStudy (anonymous):

Wow, i surprised my own self LOL

OpenStudy (anonymous):

But we still need to find y. It is helpful when working these problems to put a star by our first equation we simplified which was \[y=2x-3\]This will help us go back and find it faster.

OpenStudy (anonymous):

Now that we know the value for x what do you think we have to do to solve for y??

OpenStudy (anonymous):

i'm not exactly sure, but I am guessing we have to plug 5 in somewhere. near the x i'm thinking.

OpenStudy (anonymous):

because x=5 we can plug it into the equation \[y=2x-3\] and solve

OpenStudy (anonymous):

What did you get for the answer??

OpenStudy (anonymous):

y=7?

OpenStudy (anonymous):

Correcttt Againnnn!!!!

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

So, to put those in ordered pair. x=5 y=7

OpenStudy (anonymous):

Now what was the last part of the question? Do you have to write the answer as an ordered pair?

OpenStudy (anonymous):

(5,7)?

OpenStudy (anonymous):

(x, y) that is the format of an ordered pair

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

Do you understand algebra steps a little bit better??

OpenStudy (anonymous):

THANK YOU!!!!! Thanks for taking the time to TEACH me and not blurt out the answer. :)

OpenStudy (anonymous):

algebra is just a bunch of rearranging equations to look prettier but essentially the original equation is deep inside still. Like a girl that wears too much make up

OpenStudy (anonymous):

You're welcome

OpenStudy (anonymous):

Stevns, while we are here if I post a problem in this box will you check my answer?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

The perimeter of a rectangle is 38 ft. The length is 3 ft longer than the width. Find the dimensions. Write a system of linear equations and solve the resulting system. Let x be the length and y be the width. Write the first equation. 2x+2y=(38) Write the second equation. X=y+(8) What is the length? ( 11 _) ft What is the width? ( 8 ) ft Check the answers in the ( ) for me please :)

OpenStudy (anonymous):

I don't want to get any of it wrong LOL

OpenStudy (anonymous):

or maybe x=y+3, not 8. now i'm doubting myself lol

OpenStudy (anonymous):

I would use the letters w and l not to get confused. The numbers are correct. How did you get 8 for the second equation. I was thinking \[l=w+3\]Because the length is 3 more than the width

OpenStudy (anonymous):

so three is right not 8. i am glad i asked. i knew something looked wrong lol

OpenStudy (anonymous):

if you plug l= w + 8 into the equation (like the last problem we worked) it doesnt work out... Ya good thing you asked. You can always work the problem backwards when you're done to check your work.

OpenStudy (anonymous):

Can you check one last one? I PROMISE! lol

OpenStudy (anonymous):

omg. lol

OpenStudy (anonymous):

A tugboat goes 120 miles upstream in 20 hours. The return trip downstream takes 10 hours. Find the speed of the tugboat without a current and the speed of the current. The speed of the tugboat is ( 9 ) mph, and the speed of the current is ( 3 ) mph.

OpenStudy (anonymous):

I don't want to fail and your a good teacher/checker! LOL :)

OpenStudy (anonymous):

speed is distance over time. Upstream the boat is fighting the current and downstream it is with the current so I would setup my equations like this. Let b stand for the speed of the boat and c stand for the speed of the current.\[b \frac{ mi }{ hr }-c \frac{ mi }{ hr }=\frac{ 120mi }{ 20hr }\]\[b \frac{ mi }{ hr }+c \frac{ mi }{ hr }=\frac{ 120mi }{ 10hr }\]

OpenStudy (anonymous):

so simpler\[b-c=6;b+c=12\]since all the units are the same we can add the numbers without worry.

OpenStudy (anonymous):

Which I get 6 and 12

OpenStudy (anonymous):

solve for one variable and substitute and I get 9 for boat and 3 for current like YOOOUUUU

OpenStudy (anonymous):

YOU rule! i WISH i COULD give you a 1,000,000 medals. :) HUGS!

OpenStudy (anonymous):

You're welcome. I'm just watching football no big deal.

OpenStudy (anonymous):

LOL! I hope your team wins!

OpenStudy (anonymous):

THANK YOU!

OpenStudy (anonymous):

what class r u taking

OpenStudy (anonymous):

Algebra I and i'm not the brightest student in the class. LOL (as you can see why i'm here) :)

OpenStudy (anonymous):

everybody learns at their own pace. I am usually lost til a switch is flipped. You just need to find that switch. You'll be fine. You'll get out of it what you put in, sounds like you're trying hard.

OpenStudy (anonymous):

Good Luck!

OpenStudy (anonymous):

Thanks so much. :) --- See you around i'm sure. Have a goodnight.

OpenStudy (anonymous):

hey, I need help, just tell me the because I'm on homebound and I just wanna get it done... I hate math with a passion. here's the equations. Complete the following ordered pairs for the equation. 6x + -5y = 20 (_, -4) (_, 2)